Use Cramer’s rule to solve for y without solving for x, z and w in the system 2w + x + y + z = 3 −8w − 7x − 3y + 5z = −3 w + 4x + y + z = 6 w + 3x + 7y − z = 1 


1
Expert's answer
2021-05-25T08:18:46-0400

Solution :-

we have given the equations


"\begin{matrix}
 2w + x + y + z = 3 \\
 −8w − 7x − 3y + 5z = −3 \\
 w + 4x + y + z = 6 \\
w + 3x + 7y − z = 1
\end{matrix}"

And now write the matrix for the system and check that the determinant is not equal zero.



"A=\begin{bmatrix}
 2 & 1 & 1 & 1 \\
 -8 & -7 & -3 & 5 \\
1 & 4 & 1 & 1 \\
1 & 3 & 7 & -1 \\
\end{bmatrix}""\det A=\begin{vmatrix}
 2 & 1 & 1 & 1 \\
 -8 & -7 & -3 & 5 \\
 1 & 4 & 1 & 1 \\
 1 & 3 & 7 & -1 \\
\end{vmatrix}""=2\begin{vmatrix}
 -7 & -3 & 5 \\
 4 & 1 & 1 \\
 3 & 7 & -1 \\
\end{vmatrix}-\begin{vmatrix}
 -8 & -3 & 5 \\
 1 & 1 & 1 \\
 1 & 7 & -1 \\
\end{vmatrix}""+\begin{vmatrix}
 -8 & -7 & 5 \\
 1 & 4 & 1 \\
 1 & 3 & -1 \\
\end{vmatrix}-\begin{vmatrix}
 -8 & -7 & -3 \\
 1 & 4 & 1 \\
 1 & 3 & 7 \\
\end{vmatrix}""\begin{vmatrix}
 -7 & -3 & 5 \\
 4 & 1 & 1 \\
 3 & 7 & -1 \\
\end{vmatrix}=-4\begin{vmatrix}
 -3 & 5 \\
 7 & -1 \\
\end{vmatrix}+\begin{vmatrix}
 -7 & 5 \\
 3 & -1 \\
\end{vmatrix}""-\begin{vmatrix}
 -7 & -3 \\
 3 &7 \\
\end{vmatrix}=-4(-32)+(-8)-(-40)=160""\begin{vmatrix}
 -8 & -3 & 5 \\
 1 & 1 & 1 \\
 1 & 7 & -1 \\
\end{vmatrix}=-\begin{vmatrix}
 -3 & 5 \\
 7 & -1 \\
\end{vmatrix}+\begin{vmatrix}
 -8 & 5 \\
 1 & -1 \\
\end{vmatrix}""-\begin{vmatrix}
 -8 & -3 \\
 1 &7 \\
\end{vmatrix}=-(-32)+(3)-(-53)=88"





"\begin{vmatrix}
 -8 & -7 & 5 \\
 1 & 4 & 1 \\
 1 & 3 & -1 \\
\end{vmatrix}=-8\begin{vmatrix}
 4 & 1 \\
 3 & -1 \\
\end{vmatrix}-\begin{vmatrix}
 -7 & 5 \\
 3 & -1\\
\end{vmatrix}""+\begin{vmatrix}
 -7 & 5 \\
 4 &1 \\
\end{vmatrix}=-8(-7)-(-8)+(-27)=37"





"\begin{vmatrix}
 -8 & -7 & -3 \\
 1 & 4 & 1 \\
 1 & 3 & 7 \\
\end{vmatrix}=-8\begin{vmatrix}
 4 & 1 \\
 3 & 7 \\
\end{vmatrix}-\begin{vmatrix}
 -7 & -3 \\
 3 & 7 \\
\end{vmatrix}""+\begin{vmatrix}
 -7 & -3\\
 4 & 1 \\
\end{vmatrix}=-8(25)-(-40)+(5)=-155""\det A=2(160)-88+37-(-155)""\det A=424\not=0"




"B=\begin{bmatrix}
 2 & 1 & 3 & 1 \\
 -8 & -7 & -3 & 5 \\
1 & 4 & 6 & 1 \\
1 & 3 & 1 & -1 \\
\end{bmatrix}""\det B=\begin{vmatrix}
 2 & 1 & 3 & 1 \\
 -8 & -7 & -3 & 5 \\
 1 & 4 & 6 & 1 \\
 1 & 3 & 1 & -1 \\
\end{vmatrix}""=2\begin{vmatrix}
 -7 & -3 & 5 \\
 4 & 6 & 1 \\
 3 & 1 & -1 \\
\end{vmatrix}-\begin{vmatrix}
 -8 & -3 & 5 \\
 1 & 6 & 1 \\
 1 & 1 & -1 \\
\end{vmatrix}""+3\begin{vmatrix}
 -8 & -7 & 5 \\
 1 & 4 & 1 \\
 1 & 3 & -1 \\
\end{vmatrix}-\begin{vmatrix}
 -8 & -7 & -3 \\
 1 & 4 & 6 \\
 1 & 3 & 1 \\
\end{vmatrix}"





"\begin{vmatrix}
 -7 & -3 & 5 \\
 4 & 6 & 1 \\
 3 & 1 & -1 \\
\end{vmatrix}=-4\begin{vmatrix}
 -3 & 5 \\
 1 & -1 \\
\end{vmatrix}+6\begin{vmatrix}
 -7 & 5 \\
 3 & -1 \\
\end{vmatrix}""-\begin{vmatrix}
 -7 & -3 \\
 3 &1 \\
\end{vmatrix}=-4(-2)+6(-8)-(2)=-42"





"\begin{vmatrix}
 -8 & -3 & 5 \\
 1 & 6 & 1 \\
 1 & 1 & -1 \\
\end{vmatrix}=\begin{vmatrix}
 -3 & 5 \\
 6 & 1 \\
\end{vmatrix}-\begin{vmatrix}
 -8 & 5 \\
 1 & 1 \\
\end{vmatrix}""-\begin{vmatrix}
 -8 & -3 \\
 1 & 6 \\
\end{vmatrix}=-33-(-13)-(-45)=25"





"\begin{vmatrix}
 -8 & -7 & 5 \\
 1 & 4 & 1 \\
 1 & 3 & -1 \\
\end{vmatrix}=-8\begin{vmatrix}
 4 & 1 \\
 3 & -1 \\
\end{vmatrix}-\begin{vmatrix}
 -7 & 5 \\
 3 & -1\\
\end{vmatrix}""+\begin{vmatrix}
 -7 & 5 \\
 4 &1 \\
\end{vmatrix}=-8(-7)-(-8)+(-27)=37"





"\begin{vmatrix}
 -8 & -7 & -3 \\
 1 & 4 & 6 \\
 1 & 3 & 1 \\
\end{vmatrix}=-8\begin{vmatrix}
 4 & 6 \\
 3 & 1 \\
\end{vmatrix}-\begin{vmatrix}
 -7 & -3 \\
 3 & 1 \\
\end{vmatrix}""+\begin{vmatrix}
 -7 & -3\\
 4 & 6 \\
\end{vmatrix}=-8(-14)-(2)-(30)=80"





"\det B=2(-42)-25+3(37)-80=-78""y=\dfrac{\det B}{\det A}=\dfrac{-78}{424}=-\dfrac{39}{212}"



"\boxed{y=-\dfrac{39}{212}}answer"